OpenStudy (anonymous):

Prove that (x + y)^2 >= 4xy

OpenStudy (kc_kennylau):

\[\large\hspace{11pt}(x+y)^2-4xy\\=x^2+2xy+y^2-4xy\\=x^2-2xy+y^2\\=(x-y)^2\\\ge0\\\because(x+y)^2-4xy\ge0\\\therefore(x+y)^2\ge4xy\]

OpenStudy (anonymous):

Thanks!

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (kc_kennylau):

I love your face :D

OpenStudy (anonymous):

Nice smile, isn't it? :D

OpenStudy (kc_kennylau):

yep xD

OpenStudy (kc_kennylau):

have a great great great day :D bye :D

OpenStudy (anonymous):

Have a good one yourself! :D